OpenStudy (anonymous):

Help!!!!!!!!!! Find the inverse of f(x)= x^2 over x^2 +1 and show that f(f^-1(x))=x

OpenStudy (anonymous):

unlikely that this function has an inverse function, because it is not one to one on the other hand, you can get a choice by solving \[x=\frac{y^2}{y^2+1}\] for \(y\)

OpenStudy (anonymous):

damn typo start with \[x(y^2+1)=y^2\] then \[xy^2+x=y^2\] so \[xy^2-y^2=-x\] factor to get \[(x-1)y^2=-x\] so \[y^2=\frac{-x}{x-1}=\frac{x}{1-x}\]

OpenStudy (anonymous):

but that is \(y^2\) not \(y\) so you need \[y=\pm\sqrt{\frac{x}{1-x}}\] the plus or minus part means it is not a function